Anonymous1babes:

Drey makes $10 an hour plus $15 an hour for every hour of overtime. Overtime hours are any hours more than 40 hours for the week. Part A: Create an equation that shows the amount of wages earned, W, for working x hours in a week when there is no overtime. (3 points) Part B: Create an equation that shows the amount of wages earned, S, for working y hours of overtime. Hint: Remember to include in the equation the amount earned from working 40 hours. (3 points) Part C: Drey earned $490 in 1 week. How many hours (regular plus overtime) did he work? Show your work. (4 points) (10 points)

juju2021:

W= S$10 per hour and Y $15 per over time hour 10s+15y

Anonymous1babes:

thanks

juju2021:

np keep them coming

Anonymous1babes:

haha u enjoying this

juju2021:

ngl yes i am happy to help poeple

Anonymous1babes:

ahhh thats nice

juju2021:

keep them coming

Anonymous1babes:

Solve for x: |2x + 12| = 18 (5 points) a x = 3, x = −3 b x = 3, x = −15 c x = 15, x = −15 d x = −3, x = 15

juju2021:

b

Anonymous1babes:

The number of State Parks in California is 36 more than one third the number of State Parks in Florida. There are 85 State Parks in California. Which of the following equations would you use to find the number of State Parks in Florida? Note: f represents the number of state parks in Florida. (5 points) a 3(f − 36) = 85 b 1 over 3 f + 36 = 85 c Quantity f minus 36 all over 3 equals 85 d Quantity f plus 36 all over 3 equals 85

juju2021:

make a new one and do it

IhelpUuHELPme:

@juju2021 wrote:
b
why is it b? dude i literally told you bout the direct answers thing -.-

juju2021:

i will stop